OpenStudy (anonymous):

write the equation of the line passing through (e^pi, square root of 2) with a slope of pi/6

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

@welshfella

OpenStudy (welshfella):

use the point slope formula y - y1 = m(x - x) where (x1,y1) is the point on the line and m = the slope

OpenStudy (welshfella):

here x1 = e^pi , y1 = sqrt 2 and m = pi/6

OpenStudy (welshfella):

so just plug theses values in

OpenStudy (anonymous):

but how do i simplify? y2-square root 2=pi/6(x2-e^pi)

OpenStudy (anonymous):

\[y-\sqrt{2}=\frac{\pi}{6}(x-e^{\pi})\] Although not necessary, a linear equation in 2 variables may be simplified into the form of \[ax+by+c=0\] Let's get rid of that 6 first by multiplying throughout by 6 \[6y-6\sqrt{2}=\pi(x-e^{\pi})\] Now we simply open the parenthesis on the right side \[6y-6\sqrt{2}=\pi x-\pi e^{\pi}\] Now it's a matter of rearranging \[\pi x-6y+6\sqrt{2}-\pi e^{\pi}=0\] This is of the form \[ax+by+c=0\] In this case, \[a=\pi\]\[b=-6\]\[c=6\sqrt{2}-\pi e^{\pi}\]